A bond that matured in 20 years has a coupon rate of 12.00% and pays interest semiannually.What is the value of this bond today if similar bonds have a 14.00% yield to maturity ?
Answer #1

A B C 1 2 Coupon rate (coupon rate/2) 6.00% per half year 3 Face value of bond, FV $ 1,000 4 No. of periods ,nper (No. of yea

Calculations-

A B 1 =129/2 per half year 1000 =20*2 half years per half year 2 Coupon rate (coupon rate/2) 3 Face value of bond, FV 4 No. o

Here,
Fv is the future value of bond
rate is the periodic interest rate
nper is no. of periods
pmt is the periodic payments
Pv is the present value of bond
type = 0 if the amount is paid at the end of the year ,1 if it is paid at the beginning.

As no method was mentioned I have solved the problem using excel function.Please upvote if the answer is helpful.In case of doubt,do comment.Thanks.